Find the equivalent acute angle of 13.5


1
Expert's answer
2022-06-07T07:47:10-0400
"13.5-4\pi\approx0.93363(rad)"

"0<13.5-4\pi<\dfrac{\pi}{2}"

The equivalent acute angle of 13.5 radian is "0.93363" radian.